OpenStudy (anonymous):

please help to find the intergral of Vdv/[g(v^2-2500)]=-the intergral of dx its a differential equation

OpenStudy (phi):

let u= v^2-2500 and du = 2 v dv so dv= du/2v

OpenStudy (anonymous):

how do i treat g

OpenStudy (phi):

I assume g is a constant?

OpenStudy (phi):

I was assuming g was the gravitational constant.

OpenStudy (anonymous):

thax

OpenStudy (anonymous):

Phi is it proper to put 1/g before the intergral sign

OpenStudy (phi):

yes